OpenStudy (anonymous):

HELP ASAP: Solve by factoring: 4x^2 + 28x - 32 = 0

OpenStudy (jdoe0001):

\(\large 4x^2 + 28x - 32 = 0\implies \begin{array}{cccllll} 4(x^2&+7x&-8)\\ &\uparrow &\uparrow \\ &8-1&8\cdot -1 \end{array}\)

OpenStudy (anonymous):

yes I got that part but you have to use the zero property

OpenStudy (jdoe0001):

so... add it away then :)

OpenStudy (jdoe0001):

what binomial did you get from the "grouped set"?

OpenStudy (jdoe0001):

grouped that parenthesized that is

OpenStudy (anonymous):

(X+8)(x-1)